# A spin on Hagedorn temperatures and string stars

### v1

@@ Abstract
We discuss the correspondence between highly excited strings and black holes in the presence of angular momentum. At fixed imaginary angular velocity $\nu$, we show that free strings exhibit a Hagedorn instability, explained—as in Atick-Witten—by a thermal-winding mode turning tachyonic. This allows us to determine the exact Hagedorn temperature $\beta_{H}(\nu)$ for bosonic, type II, and heterotic strings. Using the effective field theory for the thermal-winding mode around the $\nu$-dependent background, we find a novel ‘rotating string star’ saddle (perturbatively in the angular velocity) and study its properties. This configuration describes a self-gravitating bound state of highly excited rotating strings. As in the non-rotating case, the saddle is qualitatively shown to interpolate between the rotating strings phase and a rotating black hole. We also comment on the implications of these results for anti-de Sitter space.
@@ 1 Introduction
Over time, black holes tend to shrink as they radiate. The fate of small (highly-curved) black holes in a UV-complete theory of gravity is therefore an interesting self-consistency check Bedroya:2022twb ; Basile:2023blg ; Bedroya:2024ubj ; Bedroya:2024uva ; Herraez:2024kux ; Herraez:2025gcf ; Herraez:2025clp , which could teach us a lesson on black hole microscopy. In string theory, the nature of black holes as their horizon size reaches the string scale remains an open question Susskind:1993ws ; Maldacena:1996ky ; Horowitz:1996nw ; Peet:2000hn . One version of this question is phrased in terms of Euclidean gravity Chen:2021dsw . The Euclidean black hole saddle in asymptotically $\mathbb{R}^{d}\times S^{1}_{\beta}$ string theory [^1] At finite $g_{s}$, it is impossible to fix the asymptotic length $\beta$ Atick:1988si . We work at arbitrarily small $g_{s}\ll 1$ throughout the paper, in which the problem disappears. Another option is to embed the discussion inside thermal AdS (see Urbach:2022xzw and section 5). is well described by Einstein gravity for $\beta\gg l_{s}$, but what is its description for $\beta\sim l_{s}$? Horowitz and Polchinski Horowitz:1997jc suggested that at high temperatures and for $3\leq d\leq 5$, the saddle is a normalizable condensate of strings winding on the thermal cycle of $\mathbb{R}^{d}\times S^{1}_{\beta}$. The saddle, interpreted as a self-gravitating stringy condensate Damour:1999aw ; Khuri:1999ez , is termed ‘the Horowitz-Polchinski solution’ or a ‘string star’. Surprisingly, this saddle carries an $O(G_{N}^{-1})$ thermal entropy, and qualitatively agrees with black hole thermodynamics at lower temperatures. [^2] In Chen:2021dsw a worldsheet index argument was given against a smooth transition between the two descriptions in type II string theory. To the best of our understanding, the argument carries over to our setting ($\nu\neq 0$) as well. Recently, there has been a renewed interest in string stars, extending the original setting Chen:2021dsw ; Mazel:2024alu to higher dimensions $d>5$ Balthazar:2022szl ; Balthazar:2022hno ; Bedroya:2024igb ; Chu:2025boe ; Chu:2025kzl , and other background geometries Urbach:2022xzw ; Urbach:2023npi ; Halder:2023nlp ; Agia:2023skp ; Ishibashi:2025qwn ; Chu:2024ggi ; Emparan:2024mbp ; Chu:2025fko .
@@
To gather more evidence for this proposal, it is natural to consider extra parameters to turn on, both at the string star and the black hole sides of the correspondence. In this work, we study the correspondence in the presence of angular momenta. [^3] Rotating black holes at low temperatures (near-extremal limit) have enjoyed renewed interest recently Sen:2012cj ; Moitra:2019bub ; Kapec:2023ruw ; Rakic:2023vhv ; Maulik:2024dwq due to their strong quantum effects at the near-horizon Bardeen:1999px region. Let $J$ be a rotation generator in a two-dimensional plane of $\mathbb{R}^{d}$. To add angular momenta, we consider an ensemble of fixed temperature $\beta$, and fixed imaginary ‘angular velocity’ potential $\nu$, conjugate to $J$, with the partition function
@@
We begin in section 2 by reviewing the thermodynamic properties of highly-excited rotating strings. [^4] For previous literature on this subject, see Russo:1994ev ; Matsuo:2009sx . Recently Ceplak:2023afb ; Ceplak:2024dxm studied aspects of the correspondence between free strings and black holes with angular momenta. While some of our work overlaps with theirs, there are important differences. Most importantly, they compared black holes with free strings, whereas we compare them to the string star. Secondly, all the objects we consider in this work (strings, string stars, or black holes) will possess a ‘small’ amount of angular momentum, and will remain away from extremality. More abstractly, the process that interpolates between strings and black holes in Ceplak:2023afb ; Ceplak:2024dxm was an adiabatic (fixed entropy) process of varying the string coupling. We follow Chen:2021dsw , and change the temperature (and possibly the angular velocity) at fixed string coupling (fixed theory). We believe this difference is not consequential. By studying free strings in the $\beta,\nu$ ensemble (1), we find it shares an important property with the canonical (non-rotating $\nu=0$) ensemble: it has a Hagedorn instability! Explicitly, we find that due to the high energy spectrum of free rotating strings, $Z(\beta,\nu)$ diverges at temperatures above a $\nu$-dependent Hagedorn temperature $\beta<\beta_{H}$. [^5] While we were not aware of this work until the final stages of writing,  Mertens:2014nca is the only previous work we found that presents $\beta_{H}(\nu)$, in the context of the pure NS-NS bosonic AdS3 background. Defining $R_{H}=\beta_{H}/(2\pi)$, we find (for $|\nu|<1$)
@@
We give two alternative ways to find (2). One is to directly check the Hagedorn growth of the partition function due to high-energy excitations. A second, more instructive way for our purposes is to follow Atick and Witten Atick:1988si and interpret the Hagedorn instability as a tachyonic instability of a thermal winding mode in $\mathbb{R}^{d}\times S^{1}$. Studying the winding sector of the twisted $\mathbb{R}^{d}\times S^{1}$ worldsheet theory, we were able to derive (2). As we explain, it is also possible to derive (2) perturbatively in $\nu$ from a target space analysis. Using a Kaluza-Klein (KK) reduction, we find a novel ($\nu$ dependent) effective field theory (EFT) for the thermal winding mode on $\mathbb{R}^{d}$. In this way, the Hagedorn temperature (2) is the temperature at which a winding mode turns massless.
@@
In section 3, we follow Horowitz:1997jc and employ this $d$-dimensional EFT of the winding mode coupled to gravity to find a normalizable winding condensate, or a rotating string star. [^6] Recently, Santos:2024ycg presented an attempt to find a rotating winding condensate. To our understanding, their underlying assumption is significantly different from ours. Most importantly, the asymptotic geometry of their solution was $\mathbb{R}^{d}\times S^{1}_{\beta}$ with no asymptotic angular velocity ($\nu=0$). As a result, the authors also introduce an intrinsic spin to the winding profile, and work only with heterotic string theory (in which the winding mode current sources the KK vector). Due to the asymptotic angular velocity, our equations of motion differ at leading order. Our winding profile has no intrinsic spin and is invariant to the remaining $SO(2)\times SO(d-2)$ rotation symmetry. While we consider, for simplicity, only bosonic and type II string star configurations, we believe a solution for heterotic strings exists and will share similar properties. To find the solution, we employ perturbation theory in $\nu\ll 1$, which to leading order amounts to a rigid-body approximation. By analytically continuing the solution back to real angular velocity ($2\pi\nu i=\beta\Omega$), we write down the thermodynamic properties of the rotating string star saddle. The rotating string star shares many properties with the non-rotating solution, such as $O(G_{N}^{-1})$ entropy. The winding profile is invariant to the remaining $SO(2)\times SO(d-2)$ rotation symmetry group and thus has no intrinsic spin. Notably, this solution still carries a non-zero $O(G_{N}^{-1})$ angular momentum, due to the non-trivial $\nu$ dependence, similar to the rotating black hole. In section 4, we compare these properties to rotating black holes and to free strings. As in the non-rotating case, we find qualitative agreement between both in their intermediate parameter regimes.
@@
Finally, in section 5 we follow Urbach:2022xzw ; Urbach:2023npi and discuss the extension of our results to anti-de Sitter (AdS) space. We compute the leading behavior of the AdS Hagedorn temperature $R_{H}(\nu)$ in $l_{s}/l_{ads}$, and comment on the expected properties of the rotating AdS string star.
@@ 2 The thermodynamics of highly-excited rotating strings
We begin by discussing the thermodynamic properties of free rotating strings at high energies. We will consider either bosonic, type II, or heterotic (critical) string theories in flat space, all at zero string coupling $g_{s}=0$. In this limit, the canonical partition function is well-defined Atick:1988si and the free string phase is dominant for arbitrarily high energies. We denote the number of spatial directions by $d$ ($d=25$ in the bosonic case, and $d=9$ for the superstring) and the plane of rotation by coordinates $X^{1},X^{2}$. Denoting the worldsheet spatial coordinate by $\sigma\sim\sigma+l$, the worldsheet angular momentum generator is given by (for the superstring, there is a second fermionic term)
@@ 2.1 The microcanonical ensemble
Fixing the mass $M$ and the angular momentum $J$, we would like to count the number of microstates $S(M,J)$ for high energies $\alpha^{\prime}M^{2}\gg 1$. For a recent review about free rotating strings in the microcanonical ensemble, see Ceplak:2023afb ; Ceplak:2024dxm . In the high-energy limit, the string theory partition function is dominated by the single string partition function Mertens:2015ola . Therefore, without loss of generality, we will focus on the single-string entropy $S(M,J)$. For a given mass $M$, the maximal possible amount of angular momentum is constructed by appropriately exciting the $X^{1}$, $X^{2}$ oscillators. The mass shell in this extreme case is given by $\frac{1}{2}\alpha^{\prime}M^{2}=J+c$, where $c$ is a theory-dependent fixed number. Therefore, at high energies $\alpha^{\prime}M^{2}\gg 1$, the angular momentum is bounded by the Regge trajectory
@@
as in the standard single-string entropy $S(M)$ Huang:1970iq ; Gross:1985fr ; Bowick:1985az ; Atick:1988si . The linear growth of the entropy signals a limiting temperature for the system $T_{H}=1/\beta_{H}$, termed the Hagedorn temperature Hagedorn:1965st . It is useful to write down the temperature in terms of the “thermal radius” $R_{H}=\beta_{H}/(2\pi)$,
@@
Based on the aforementioned bound (4), a more natural scaling is to keep $1/2<J/(\alpha^{\prime}M^{2})<1/2$ fixed as we take the high-energy limit. As we show in appendix A.2 (see also Matsuo:2009sx ), the high energy behavior is
@@
A third possible high-energy ($l_{s}M\gg 1$) limit is to keep $J/(l_{s}M)$ fixed. We can further expand in both large $l_{s}M\gg 1$ and small $J/(l_{s}M)\ll 1$. For a derivation of the analysis, see appendix A.3 and Russo:1994ev for more details. The entropy has the double expansion
@@
with $\alpha_{1},\alpha_{2}$ theory-dependent constants [^7] In case where some of the dimensions are compact, this formula needs to be corrected Mertens:2015ola .
@@
Following Ceplak:2024dxm , we would also like to estimate the shape of the rotating string. As an estimate for the typical sizes of the string in the rotation plane ($L_{||}$) and in the directions orthogonal to it ($L_{\perp}$), one can calculate the variance of the corresponding coordinate $\langle(X^{i})^{2}\rangle$. In appendix A.3, we analyze the case of bosonic strings, which gives
@@ 2.2 The grand canonical ensemble
This ensemble is especially familiar in the context of AdS/CFT Hawking:1998kw ; Hawking:1999dp . Unfortunately, in flat space, fixing the angular velocity without fixing the energy is generically unstable. The reason is that it is always possible to excite the system with low energy but arbitrarily high angular momentum $\Delta J$, so that the Boltzmann factor will be enhanced by $\sim\exp(+\Omega\ \Delta J)$ Witten:2021nzp ; Kim:2023sig . Allowing arbitrarily high $\Delta J$ will result in a diverging partition function. [^8] In other words, the multi-string partition function is divergent. As a matter of fact, the single-string partition function is also divergent for a somewhat different reason. As we explained above, for high enough energies, the system is dominated by maximally rotating strings. While the Boltzmann factor is $\exp(-\beta M)$, the potential enhances it by $\sim\exp(\beta\Omega M^{2}/2)$. Therefore, the sum over $M$ results in a diverging partition function. See appendix B for a classical intuition for this result.
@@
While it might look like an odd thing to do, note that, for example, in AdS3 Maldacena:1998bw ; Maloney:2007ud this construction is natural, and dual to the familiar CFT2 torus partition function. Similarly, it is in this ensemble that the rotating black hole solution is real (in Euclidean signature). On the other hand, this ensemble is harder to interpret statistically, as it gives complex probabilities for microstates.
@@
In the remainder of this section, we will show that for bosonic strings. We refer the reader to appendix A for a derivation also for type II and heterotic strings. Our reasoning will follow an analysis of Russo:1994ev ; Matsuo:2009sx , or more recently by Ceplak:2024dxm . We begin by working in light-cone quantization, and computing the worldsheet partition function with fixed worldsheet temperature $1/b$ and imaginary (target space) angular velocity $\nu$. The (single chirality) result for the $d$-dimensional bosonic string is ($x=e^{-b}$)Matsuo:2009sx
@@
With $\eta$ the Dedekind function and $\vartheta_{11}$ the Jacobi $\theta$ function Polchinski:1998rq . Using the modular properties of the expression, the small $b$ behavior is

### v2

@@ Abstract
We discuss the correspondence between highly excited strings and black holes in the presence of angular momentum. At fixed imaginary angular velocity $\nu$, we show that free strings exhibit a Hagedorn instability due to a thermal-winding mode turning tachyonic. This allows us to determine the exact Hagedorn temperature $\beta_{H}(\nu)$ for bosonic, type II, and heterotic strings. Using the effective field theory for the thermal-winding mode around the $\nu$-dependent background, we find a novel ‘rotating string star’ saddle (perturbatively in the angular velocity) and study its properties. This configuration describes a self-gravitating bound state of highly excited rotating strings. As in the non-rotating case, the saddle is qualitatively shown to interpolate between the rotating strings phase and a rotating black hole. We also comment on the implications of these results for anti-de Sitter space.
@@ 1 Introduction
Over time, black holes tend to shrink as they radiate. The fate of small (highly-curved) black holes in a UV-complete theory of gravity is therefore an interesting self-consistency check Bedroya (2022); Basile et al. (2024); Bedroya et al. (2025, 2024); Herráez et al. (2025b, c, a), which could teach us a lesson on black hole microscopy. In string theory, the nature of black holes as their horizon size reaches the string scale remains an open question Susskind (1993); Maldacena (1996); Horowitz and Polchinski (1997); Peet (2000). One version of this question is phrased in terms of Euclidean gravity Chen et al. (2023). The Euclidean black hole saddle in asymptotically $\mathbb{R}^{d}\times S^{1}_{\beta}$ string theory [^1] At finite $g_{s}$, it is impossible to fix the asymptotic length $\beta$ Atick and Witten (1988). We work at arbitrarily small $g_{s}\ll 1$ throughout the paper, in which the problem disappears. Another option is to embed the discussion inside thermal AdS (see Urbach (2022) and section 5). is well described by Einstein gravity for $\beta\gg l_{s}$, but what is its description for $\beta\sim l_{s}$? Horowitz and Polchinski Horowitz and Polchinski (1998) suggested that at high temperatures and for $3\leq d\leq 5$, the saddle is a normalizable condensate of strings winding on the thermal cycle of $\mathbb{R}^{d}\times S^{1}_{\beta}$. The saddle, interpreted as a self-gravitating stringy condensate Damour and Veneziano (2000); Khuri (1999), is termed ‘the Horowitz-Polchinski solution’ or a ‘string star’. Surprisingly, this saddle carries an $O(G_{N}^{-1})$ thermal entropy, and qualitatively agrees with black hole thermodynamics at lower temperatures. [^2] In Chen et al. (2023) a worldsheet index argument was given against a smooth transition between the two descriptions in type II string theory. To the best of our understanding, the argument carries over to our setting ($\nu\neq 0$) as well. Recently, there has been a renewed interest in string stars, extending the original setting Chen et al. (2023); Mazel et al. (2024) to higher dimensions $d>5$ Balthazar et al. (2022, 2024); Bedroya and Wu (2025); Chu and Kutasov (2025a, b), and other background geometries Urbach (2022, 2023); Halder and Jafferis (2024); Agia and Jafferis (2023); Ishibashi et al. (2025); Chu (2025a); Emparan et al. (2025); Chu (2025b).
@@
To gather more evidence for this proposal, it is natural to consider extra parameters to turn on, both at the string star and the black hole sides of the correspondence. In this work, we study the correspondence in the presence of angular momenta. [^3] Rotating black holes at low temperatures (near-extremal limit) have enjoyed renewed interest recently Sen (2012); Moitra et al. (2019); Kapec et al. (2024); Rakic et al. (2024); Maulik et al. (2024) due to their strong quantum effects at the near-horizon Bardeen and Horowitz (1999) region. Let $J$ be a rotation generator in a two-dimensional plane of $\mathbb{R}^{d}$. To add angular momenta, we consider an ensemble of fixed temperature $\beta$, and fixed imaginary ‘angular velocity’ potential $\nu$, conjugate to $J$, with the partition function
@@
We begin in section 2 by reviewing the thermodynamic properties of highly-excited rotating strings. [^4] For previous literature on this subject, see Russo and Susskind (1995); Matsuo (2010). Recently Čeplak et al. (2023, 2024) studied aspects of the correspondence between free strings and black holes with angular momenta. While some of our work overlaps with theirs, there are important differences. Most importantly, they compared black holes with free strings, whereas we compare them to the string star. Secondly, all the objects we consider in this work (strings, string stars, or black holes) will possess a ‘small’ amount of angular momentum, and will remain away from extremality. More abstractly, the process that interpolates between strings and black holes in Čeplak et al. (2023, 2024) was an adiabatic (fixed entropy) process of varying the string coupling. We follow Chen et al. (2023), and change the temperature (and possibly the angular velocity) at fixed string coupling (fixed theory). We believe this difference is not consequential. By studying free strings in the $\beta,\nu$ ensemble (1), we find it shares an important property with the canonical (non-rotating $\nu=0$) ensemble: it has a Hagedorn instability! Explicitly, we find that due to the high energy spectrum of free rotating strings, $Z(\beta,\nu)$ diverges at temperatures above a $\nu$-dependent Hagedorn temperature $\beta<\beta_{H}$. [^5] While we were not aware of this work until the final stages of writing,  Mertens et al. (2014) is the only previous work we found that presents $\beta_{H}(\nu)$, in the context of the pure NS-NS bosonic AdS3 background. Defining $R_{H}=\beta_{H}/(2\pi)$, we find (for $|\nu|<1$)
@@
We give two alternative ways to find (2). One is to directly check the Hagedorn growth of the partition function due to high-energy excitations. A second, more instructive way for our purposes is to follow Sathiapalan (1987); Kogan (1987); Atick and Witten (1988) and interpret the Hagedorn instability as a tachyonic instability of a thermal winding mode in $\mathbb{R}^{d}\times S^{1}$. Studying the winding sector of the twisted $\mathbb{R}^{d}\times S^{1}$ worldsheet theory, we were able to derive (2). As we explain, it is also possible to derive (2) perturbatively in $\nu$ from a target space analysis. Using a Kaluza-Klein (KK) reduction, we find a novel ($\nu$ dependent) effective field theory (EFT) for the thermal winding mode on $\mathbb{R}^{d}$. In this way, the Hagedorn temperature (2) is the temperature at which a winding mode turns massless.
@@
In section 3, we follow Horowitz and Polchinski (1998) and employ this $d$-dimensional EFT of the winding mode coupled to gravity to find a normalizable winding condensate, or a rotating string star. [^6] Recently, Santos and Zigdon (2024) presented an attempt to find a rotating winding condensate. To our understanding, their underlying assumption is significantly different from ours. Most importantly, the asymptotic geometry of their solution was $\mathbb{R}^{d}\times S^{1}_{\beta}$ with no asymptotic angular velocity ($\nu=0$). As a result, the authors also introduce an intrinsic spin to the winding profile, and work only with heterotic string theory (in which the winding mode current sources the KK vector). Due to the asymptotic angular velocity, our equations of motion differ at leading order. Our winding profile has no intrinsic spin and is invariant to the remaining $SO(2)\times SO(d-2)$ rotation symmetry. While we consider, for simplicity, only bosonic and type II string star configurations, we believe a solution for heterotic strings exists and will share similar properties. To find the solution, we employ perturbation theory in $\nu\ll 1$, which to leading order amounts to a rigid-body approximation. By analytically continuing the solution back to real angular velocity ($2\pi\nu i=\beta\Omega$), we write down the thermodynamic properties of the rotating string star saddle. The rotating string star shares many properties with the non-rotating solution, such as $O(G_{N}^{-1})$ entropy. The winding profile is invariant to the remaining $SO(2)\times SO(d-2)$ rotation symmetry group and thus has no intrinsic spin. Notably, this solution still carries a non-zero $O(G_{N}^{-1})$ angular momentum, due to the non-trivial $\nu$ dependence, similar to the rotating black hole. In section 4, we compare these properties to rotating black holes and to free strings. As in the non-rotating case, we find qualitative agreement between both in their intermediate parameter regimes.
@@
Finally, in section 5 we follow Urbach (2022, 2023) and discuss the extension of our results to anti-de Sitter (AdS) space. We compute the leading behavior of the AdS Hagedorn temperature $R_{H}(\nu)$ in $l_{s}/l_{ads}$, and comment on the expected properties of the rotating AdS string star.
@@ 2 The thermodynamics of highly-excited rotating strings
We begin by discussing the thermodynamic properties of free rotating strings at high energies. We will consider either bosonic, type II, or heterotic (critical) string theories in flat space, all at zero string coupling $g_{s}=0$. In this limit, the canonical partition function is well-defined Atick and Witten (1988) and the free string phase is dominant for arbitrarily high energies. We denote the number of spatial directions by $d$ ($d=25$ in the bosonic case, and $d=9$ for the superstring) and the plane of rotation by coordinates $X^{1},X^{2}$. Denoting the worldsheet spatial coordinate by $\sigma\sim\sigma+l$, the worldsheet angular momentum generator is given by (for the superstring, there is a second fermionic term)
@@ 2.1 The microcanonical ensemble
Fixing the mass $M$ and the angular momentum $J$, we would like to count the number of microstates $S(M,J)$ for high energies $\alpha^{\prime}M^{2}\gg 1$. For a recent review about free rotating strings in the microcanonical ensemble, see Čeplak et al. (2023, 2024). In the high-energy limit, the string theory partition function is dominated by the single string partition function Mertens (2015). Therefore, without loss of generality, we will focus on the single-string entropy $S(M,J)$. For a given mass $M$, the maximal possible amount of angular momentum is constructed by appropriately exciting the $X^{1}$, $X^{2}$ oscillators. The mass shell in this extreme case is given by $\frac{1}{2}\alpha^{\prime}M^{2}=J+c$, where $c$ is a theory-dependent fixed number. Therefore, at high energies $\alpha^{\prime}M^{2}\gg 1$, the angular momentum is bounded by the Regge trajectory
@@
as in the standard single-string entropy $S(M)$ Huang and Weinberg (1970); Gross et al. (1985); Bowick and Wijewardhana (1985); Atick and Witten (1988). The linear growth of the entropy signals a limiting temperature for the system $T_{H}=1/\beta_{H}$, termed the Hagedorn temperature Hagedorn (1965). It is useful to write down the temperature in terms of the “thermal radius” $R_{H}=\beta_{H}/(2\pi)$,
@@
Based on the aforementioned bound (4), a more natural scaling is to keep $1/2<J/(\alpha^{\prime}M^{2})<1/2$ fixed as we take the high-energy limit. As we show in appendix A.2 (see also Matsuo (2010)), the high energy behavior is
@@
A third possible high-energy ($l_{s}M\gg 1$) limit is to keep $J/(l_{s}M)$ fixed. We can further expand in both large $l_{s}M\gg 1$ and small $J/(l_{s}M)\ll 1$. For a derivation of the analysis, see appendix A.3 and Russo and Susskind (1995) for more details. The entropy has the double expansion
@@
with $\alpha_{1},\alpha_{2}$ theory-dependent constants [^7] In case where some of the dimensions are compact, this formula needs to be corrected Mertens (2015).
@@
Following Čeplak et al. (2024), we would also like to estimate the shape of the rotating string. As an estimate for the typical sizes of the string in the rotation plane ($L_{||}$) and in the directions orthogonal to it ($L_{\perp}$), one can calculate the variance of the corresponding coordinate $\langle(X^{i})^{2}\rangle$. In appendix A.3, we analyze the case of bosonic strings, which gives
@@ 2.2 The grand canonical ensemble
This ensemble is especially familiar in the context of AdS/CFT Hawking et al. (1999); Hawking and Reall (2000). Unfortunately, in flat space, fixing the angular velocity without fixing the energy is generically unstable. The reason is that it is always possible to excite the system with low energy but arbitrarily high angular momentum $\Delta J$, so that the Boltzmann factor will be enhanced by $\sim\exp(+\Omega\ \Delta J)$ Witten (2021); Kim et al. (2023). Allowing arbitrarily high $\Delta J$ will result in a diverging partition function. [^8] In other words, the multi-string partition function is divergent. As a matter of fact, the single-string partition function is also divergent for a somewhat different reason. As we explained above, for high enough energies, the system is dominated by maximally rotating strings. While the Boltzmann factor is $\exp(-\beta M)$, the potential enhances it by $\sim\exp(\beta\Omega M^{2}/2)$. Therefore, the sum over $M$ results in a diverging partition function. See appendix B for a classical intuition for this result.
@@
While it might look like an odd thing to do, note that, for example, in AdS3 Maldacena and Strominger (1998); Maloney and Witten (2010) this construction is natural, and dual to the familiar CFT2 torus partition function. Similarly, it is in this ensemble that the rotating black hole solution is real (in Euclidean signature). On the other hand, this ensemble is harder to interpret statistically, as it gives complex probabilities for microstates.
@@
In the remainder of this section, we will show that for bosonic strings. We refer the reader to appendix A for a derivation also for type II and heterotic strings. Our reasoning will follow an analysis of Russo and Susskind (1995); Matsuo (2010), or more recently by Čeplak et al. (2024). We begin by working in light-cone quantization, and computing the worldsheet partition function with fixed worldsheet temperature $1/b$ and imaginary (target space) angular velocity $\nu$. The (single chirality) result for the $d$-dimensional bosonic string is ($x=e^{-b}$)Matsuo (2010)
@@
With $\eta$ the Dedekind function and $\vartheta_{11}$ the Jacobi $\theta$ function Polchinski (2007a). Using the modular properties of the expression, the small $b$ behavior is